# COSINE-100 Collaboration

### v2

@@ Abstract
We present results of the search for WIMP-like dark matter interaction with electrons in the NaI(Tl) crystals of the COSINE-100 experiment. The two benchmark scenarios of a heavy and a light vector boson as mediator of the interaction were studied. We found no excess events over the expected background in a data-set of 2.82 years, with a total exposure of 172.9 kg-year. The derived 90% confidence level upper limits exclude a WIMP-electron scattering cross section above 6.4 $\times$ 10-33 cm2 for a WIMP mass of 0.25 GeV, assuming a light mediator; and above 3.4 $\times$ 10-37 cm2 for a 0.4 GeV WIMP, assuming a heavy mediator, and represent the most stringent constraints for a NaI(Tl) target to date. We also briefly discuss a planned analysis using an annual modulation method below the current 0.7 keV threshold of COSINE-100, down to few photoelectrons yield.
@@ I Introduction
The existence of cold dark matter (CDM) is supported by numerous astrophysical and cosmological observations that demonstrate its gravitational interactions. All evidence is compatible with 85% of the matter in the universe being composed of DM. Nevertheless, the precise nature of DM particles remains unknown, prompting extensive research in the field of particle physics Freese (2009); Abbott et al. (2019); Bond et al. (1984); Colombi et al. (1996); Bertone and Tait (2018). One of the main DM candidates is the Weakly Interacting Massive Particle (WIMP). Due to the WIMP miracle Feng (2023) and since different theories that extend the standard model propose particles with the WIMP properties, WIMPs are rendered as a highly promising candidate Jungman et al. (1996).
@@
To date, the most studied scenario considers a weak interaction between WIMPs and nucleons. Nonetheless, because there is no compelling evidence of WIMP detection in the results from multiple direct detection experiments utilizing different detector materials, as well as the increasingly stringent constraints on the WIMP-nucleon cross section and WIMP mass Abdelhameed et al. (2019); Agnes et al. (2023a); Aalbers et al. (2023); Aprile et al. (2019), studies of alternative models are encouraged Necib et al. (2017); Sigurdson et al. (2004); Fabbrichesi et al. (2021); Peccei (2008). One such alternative involves interactions between WIMP-like cold DM and electrons Essig et al. (2016), which could be detected in modern low threshold detectors. This is particularly relevant in the context of sub-GeV WIMPs, where increased sensitivity to this interaction is anticipated. Given that the largest unexplored region pertains to sub-GeV WIMPs, the WIMP-electron scattering is a viable and appropriate scenario.
@@
Several collaborations have reported their searches for the WIMP-electron interaction, constraining its cross section Aprile et al. (2022); Essig et al. (2012); Zhang et al. (2022); Arnaud et al. (2020). As a consequence of the low threshold, silicon based experiments, such as DAMIC-M Aggarwal et al. (2025) and SENSEI Adari et al. (2025), have set the most stringent limits for lighter WIMPs within the mass region of 1 to approximately 100 MeV. In contrast, as a result of the high detector exposure, liquid noble gas experiments, such as PANDAX-4T Li et al. (2023), DarkSide-50 Agnes et al. (2023b) and XENONnT Aprile et al. (2025), have provided the strongest constraints for WIMPs with masses exceeding 100 MeV. Despite that, up to this moment, no experiment utilizing NaI(Tl) detectors has published a dedicated search for this cold DM WIMP-electron interaction model. The NEON experiment has recently published a search for light DM scattering on electrons in NaI(Tl) detectors Choi et al. (2025), probing DM masses in the keV range. However, the studied model consisted in DM particles produced by the invisible decay of dark photons generated in the core of a nuclear reactor, which only enabled the search for DM masses up to 1 MeV. Some studies have interpreted the long-lasting DAMA/LIBRA modulation results Bernabei et al. (2018, 2021) in relation to the detection of WIMP scattering on electrons Roberts et al. (2016); Roberts and Flambaum (2019). These studies have determined the values for WIMP mass and cross section with electrons related to the best-fit to DAMA modulation data. While WIMP-electron scattering alone cannot fully account for the modulated signal across the entire DAMA/LIBRA energy spectrum, it has been proposed as a potential contributor, particularly in the low-energy region near 1 keV. Also, especially with the threshold reduction of DAMA/LIBRA’s crystals to 0.75 keV, this model can only partially explain DAMA’s modulation results, with good fitting near the 1 keV energy range, but poor accordance for energies up to 6 keV. However, due to the lack of searches for the WIMP-electron scattering model in NaI(Tl) detectors, investigating it with COSINE-100 can be highly useful for confirming and expanding the restrictions on the WIMP-electron cross section.
@@
This work presents the results of the search for events generated by WIMP-like DM scattering on electrons in the COSINE-100 experiment. Excess events over the expected background were searched in a data-set of 2.82 years, using 61.3 kg of NaI(Tl), totalizing an exposure of 172.9 kg-year. A reduced analysis threshold of 8 photoelectrons, corresponding to an energy of 0.7 keV, has been achieved in COSINE-100 crystals Yu et al. (2024), thereby significantly improving the sensitivity for the detection of light WIMPs scattering on electrons. The presented study addresses two traditional cases of the WIMP-electron scattering mediated by a vector boson: a mediator that is much heavier than the electron, resulting in a contact interaction, and a mediator much lighter than the electron, resulting in a long-range interaction.
@@ II COSINE-100 experiment
Active and passive vetoes were present in the COSINE-100 experiment. Surrounding the LS there were a 3 cm thick copper shield, low activity 20 cm thick lead bricks, and 37 Eljen EJ-200 plastic scintillator panels, present as an active veto for cosmic muons. The LS acts not only as a passive and active veto, but is also crucial for measuring the activity of radioisotopes contributing to the background in the crystals. Fig. 1 shows a schematic view of the COSINE-100 experiment.
@@
Each PMT attached to the crystals had a dynode, whose signals are suitable for analyses with energies above 100 keV up to few MeVs, in which signals become saturated; and an anode, whose signals are used in low energy analyses, such as the one presented in this work. Even though all 8 crystals are used in analyses above 100 keV, 3 crystals have a poor performance for low energy signals due to low light yield and high noise and background rate. Hence, only data from the remaining 5 crystals are suitable for low energy analysis, and have been used in COSINE-100 WIMP searches Carlin et al. (2025); Yu et al. . The total effective mass of these 5 crystals is 61.3 kg.
Anode signals from the crystals’ PMTs were amplified by a factor of 30 and dynode signals amplified by a factor of 100 before processed by 500 mega sample per second (MSPS) FADC modules. As for the active vetoes, signals were read by 63.5 MSPS analog to digital converter (M64ADC) modules. Signals from the plastic scintillators and crystals were in the core of the trigger logic. If the pulse height of a signal in any crystal PMT rises above a threshold of 6 mV, a coincidence gate of 200 ns was generated waiting for another signal above threshold in the other PMT in the same crystal. When the coincidence was observed, the TCB created a global trigger and data from all FADC and M64ADC modules were stored. A global trigger could also be generated if two or more plastic panels observed coincident events. Hence, LS events could only be stored if signals from at least one crystal or two plastic panels were triggered.
Variables from the laboratory environment and detector stability parameters were monitored. Detectors event rate, behavior of parameters used for event selection, and other detector aspects related to its stability were daily monitored by the collaboration. Environment variables were also controlled through the Grafana application gra , as for example the PMTs voltage levels, relative humidity, radon air concentration and the detectors temperature. More details of the COSINE-100 experiment detectors performance, data acquisition system, and environmental monitoring are reported in Refs. Adhikari et al. (2018, 2021a); Prihtiadi et al. (2018); Adhikari et al. (2022).
@@ III.1 Overview and key expressions for direct detection
According to the momentum and energy conservations, WIMPs deposit more energy in the target when scattering on particles that have a mass similar to their own. Hence, the recent focus on searches for lighter WIMPs, especially below 10 GeV, has motivated the increase in the number of studies about WIMP-like DM interaction with electrons in the last years, since light WIMPs, with mass below few GeVs, can induce a detectable amount of events above the threshold of current detectors.
Although the standard WIMP-nucleon scattering scenario assumes a weak interaction Engel et al. (1992), there are no restrictions in the models to DM interacting simultaneously with nucleons and electrons, despite a new vector boson is commonly postulated in the latter framework Essig et al. (2016). The most frequently considered model for the interaction between WIMPs and electrons suggests a hidden sector boson to be the mediator, without restrictions to its spin and parity Roberts et al. (2016). However, in order to simplify the model and accommodate the well motivated scenario of the dark photon Fabbrichesi et al. (2021), this boson is supposed to be a vector. In this work, we consider only the scattering between WIMPs and electrons, neglecting the WIMP-nucleon interaction, in order to directly compare the analysis with reported studies from other experiments. Nevertheless, the typical momentum transfer involved in the scattering is much smaller than the sodium and iodine nuclei masses, implying that the nuclear recoil energy would be negligible. In this scenario, the Feynman diagram representing the scattering is presented in Fig. 2.
Figure 2: Feynman diagram that represents the DM-electron scattering, where $V$ is a vector boson and $\chi$ is the WIMP.
Generally, describing the kinematics of the electron ionization by a WIMP scattering is more complex than the typical WIMP-nucleon interaction, because the possible values of the momentum transfer in the interaction ($q$) can be as small as O(100 eV) for WIMP masses in the MeV range and energy depositions of O(1-10 eV), or as high as O(10 MeV) for WIMP masses in the GeV range and energy depositions of O(1-10 keVs) Lee et al. (2015); Roberts et al. (2016).
Focusing on the direct detection of electron recoils, the velocity averaged cross section for the WIMP-electron scattering in a detector target can be determined by Essig et al. (2012):
@@
where $f(v_{\chi})$ and $v_{\chi}$ are the WIMP velocity distribution and WIMP velocity in the Earth frame, respectively, $f_{ion}(E,q)$ is the ionization factor, and $q_{+}$ and $q_{-}$ are the maximum and minimum momentum transfer, respectively, given by $q_{\pm}=m_{\chi}v_{\chi}\pm\sqrt{m^{2}_{\chi}v_{\chi}^{2}-2m_{\chi}E}$, where $m_{\chi}$ is the WIMP mass, and $E$ is the energy deposition.
The DM form factor ($F_{\chi}(q)$) and the WIMP-electron scattering cross section (${\overline{\sigma}}_{e}$) are the main expressions that contain information about the interaction mediator. These are given by Roberts and Flambaum (2019):
$F_{\chi}\,(q)=\frac{\big(\frac{m_{DP}}{m_{e}}\big)^{2}+\alpha^{2}}{\big(\frac{m_{DP}}{m_{e}}\big)^{2}+(\alpha\,a_{0}\,q)^{2}}$ (2)
$\overline{\sigma}_{e}=\frac{16\,\pi\,a_{0}^{2}\,\alpha^{2}\,\alpha_{\chi}^{2}}{\Big(\big(\frac{m_{DP}}{m_{e}}\big)^{2}+\alpha^{2}\Big)^{2}}$ (3)
where $\alpha$ is the fine structure constant, $m_{e}$ is the electron mass, $a_{0}$ is the Bohr radius, written as $a_{0}=\frac{1}{m_{e}\,\alpha}$, $\alpha_{\chi}$ is the WIMP-electron coupling strength. $\overline{\sigma}_{e}$ is the interaction cross section considering a free electron whose scattering with the WIMP transferred a momentum of $a^{-1}_{0}$.
Since $\overline{\sigma}_{e}$ and $F_{\chi}(q)$ depend on the mediator mass, studies usually consider two limiting cases, such that a standard scenario is established, and a precise comparison between different results becomes simple. The first case considers a heavy mediator ($m_{DP}>>\alpha\,m_{e}$), therefore, $F_{\chi}(q)\longrightarrow 1$. The opposite case considers a light mediator ($m_{DP}<<\alpha\,m_{e}$), therefore, $F_{\chi}(q)\longrightarrow\dfrac{1}{(a_{0}\,q)^{2}}$.
The standard halo model has been assumed in this analysis:
@@
where $\vec{v}$ is the WIMP velocity in the galactic frame, $\vec{v_{E}}(t)$ is the Earth velocity in the galactic frame, $v_{0}=220$ km/s is the Sun rotation velocity with respect to the galactic center, and $v_{esc}=550$ km/s is the Milky Way escape velocity. $\vec{v_{E}}(t)$ depends on the relative velocity of Earth with respect to the Sun, and the peculiar velocity of the solar system in the galactic frame. This parametrization is in accordance with the procedure described in Ref. Freese et al. (2013). The chosen parameters values are the same considered in Ref. Roberts and Flambaum (2019).
@@
where $n^{NaI}_{a}$ is the atomic number density per kilogram of NaI, and $\rho_{DM}=0.4$ GeV/cm3 is the approximate energy density of CDM at Earth. This $\rho_{DM}$ value is the benchmark model for WIMP-electron scattering searches.
It is important to note that even though the number density of electrons in thallium is higher than in sodium and iodine, its concentration in COSINE-100 crystals is very low, and can be neglected in the analysis. Naive calculations have shown that the maximum contribution of thallium is about 1% of the total number of events from NaI.
Different descriptions are considered when determining the $f_{ion}$ parameter in WIMP-electron interaction studies, and an accurate description is essential depending on the target and energies under study, as very divergent results can be obtained Caddell et al. (2023).

### v3

@@ Abstract
We present results of the search for sub-GeV dark matter interaction with electrons in the NaI(Tl) crystals of the COSINE-100 experiment. The two benchmark scenarios of a heavy and a light vector boson as mediator of the interaction were studied. We found no excess events over the expected background in a data-set of 2.82 years, with a total exposure of 172.9 kg-year. The derived 90% confidence level upper limits exclude a DM-electron scattering cross section above 6.4 $\times$ 10-33 cm2 for a DM mass of 0.25 GeV, assuming a light mediator; and above 3.4 $\times$ 10-37 cm2 for a 0.4 GeV DM, assuming a heavy mediator, and represent the most stringent constraints for a NaI(Tl) target to date. We also briefly discuss a planned analysis using an annual modulation method below the current 0.7 keV threshold of COSINE-100, down to few photoelectrons yield.
@@ I Introduction
The existence of cold dark matter (CDM) is supported by numerous astrophysical and cosmological observations that demonstrate its gravitational interactions [1, 2, 3]. All evidence is compatible with 85% of the matter in the universe being composed of DM. Nevertheless, the precise nature of DM particles remains unknown, prompting extensive research in the field of particle physics [4, 5, 6, 7]. One of the main DM candidates is the Weakly Interacting Massive Particle (WIMP). Due to the WIMP miracle [8] and since different theories that extend the standard model propose particles with the WIMP properties, WIMPs are rendered as a highly promising candidate [9].
@@
To date, the most studied scenario considers a weak interaction between WIMPs and nucleons. Nonetheless, because there is no compelling evidence of WIMP detection in the results from multiple direct detection experiments utilizing different detector materials, as well as the increasingly stringent constraints on the WIMP-nucleon cross section and WIMP mass [10, 11, 12, 13], studies of alternative models are encouraged [14, 15, 16, 17]. One such alternative involves interactions between WIMP-like cold DM and electrons [18], which could be detected in modern low threshold detectors. This is particularly relevant in the context of sub-GeV particles, where increased sensitivity to this interaction is anticipated. Given that the largest unexplored region pertains to sub-GeV WIMP-like DM, the search for the scattering of this light DM and electrons is a viable and appropriate scenario.
@@
Several collaborations have reported their searches for the light DM-electron interaction, constraining its cross section [19, 20, 21, 22]. As a consequence of the low threshold, silicon based experiments, such as DAMIC-M [23] and SENSEI [24], have set the most stringent limits for sub-GeV DM within the mass region of 1 MeV to approximately 50 MeV. In contrast, as a result of the high detector exposure and low background, liquid noble gas experiments, such as PANDAX-4T [25], DarkSide-50 [26] and XENONnT [27], have provided the strongest constraints for light DM with masses exceeding 50 MeV. Despite that, up to this moment, no experiment utilizing NaI(Tl) detectors has published a dedicated search for this cold sub-GeV DM-electron interaction model. The NEON experiment has recently published a search for light DM scattering on electrons in NaI(Tl) detectors [28], probing DM masses in the keV range. However, the studied model consisted in DM particles produced by the invisible decay of dark photons generated in the core of a nuclear reactor, which only enabled the search for DM masses up to 1 MeV. Some studies have interpreted the long-lasting DAMA/LIBRA modulation results [29, 30] in relation to the detection of DM scattering on electrons [31, 32]. These studies have determined the values for DM mass and cross section with electrons related to the best-fit to DAMA modulation data. While light DM-electron scattering alone cannot fully account for the modulated signal across the entire DAMA/LIBRA energy spectrum, it has been proposed as a potential contributor, particularly in the low-energy region near 1 keV. Also, especially with the threshold reduction of DAMA/LIBRA’s crystals to 0.75 keV, this model can only partially explain DAMA’s modulation results, with good fitting near the 1 keV energy range, but poor accordance for energies up to 6 keV. However, due to the lack of searches for the DM-electron scattering model in NaI(Tl) detectors, investigating it with COSINE-100 can be highly useful for confirming and expanding the restrictions on the DM-electron cross section.
@@
This work presents the results of the search for events generated by sub-GeV DM scattering on electrons in the COSINE-100 experiment. Excess events over the expected background were searched in a data-set of 2.82 years, using 61.3 kg of NaI(Tl), totalizing an exposure of 172.9 kg-year. A reduced analysis threshold of 8 photoelectrons, corresponding to an energy of 0.7 keV, has been achieved in COSINE-100 crystals [33], thereby significantly improving the sensitivity for the detection of sub-GeV DM scattering on electrons. The presented study addresses two traditional cases of the DM-electron scattering mediated by a vector boson: a mediator that is much heavier than the electron, resulting in a contact interaction, and a mediator much lighter than the electron, resulting in a long-range interaction.
@@ II COSINE-100 experiment
Active and passive vetoes were present in the COSINE-100 experiment. Surrounding the LS, there was a 3 cm thick copper shield and a 20 cm thick layer of low activity lead bricks. In addition, 37 Eljen EJ-200 plastic scintillator panels provided an active veto against cosmic muons. The LS acted not only as a passive and active veto, but was also crucial for measuring the activity of radioisotopes contributing to the background in the crystals. Fig. 1 shows a schematic view of the COSINE-100 experiment.
@@
Each PMT coupled to the crystals provided both dynode and anode readouts. The dynode signals are used for analyses in the energy range above 100 keV up to a few MeVs, where saturation effects become relevant. In contrast, the anode signals were employed for low energy analyses, such as the one presented in this work. Even though all 8 crystals are used in analyses above 100 keV, 3 crystals have a poor performance for low energy signals due to low light yield and high noise and background rate. Hence, only data from the remaining 5 crystals are suitable for low energy analysis, and have been used in COSINE-100 WIMP searches [34, 35]. The total effective mass of these 5 crystals is 61.3 kg.
Anode signals from the crystals’ PMTs were amplified by a factor of 30 and dynode signals amplified by a factor of 100 before processed by 500 mega sample per second (MSPS) FADC modules. As for the active vetoes, signals were read by 62.5 MSPS analog to digital converter (M64ADC) modules. Signals from the plastic scintillators and crystals were in the core of the trigger logic. If the pulse height of a signal in any crystal PMT rises above a threshold of 6 mV, a coincidence gate of 200 ns was generated waiting for another signal above threshold in the other PMT in the same crystal. When the coincidence was observed, the TCB created a global trigger and data from all FADC and M64ADC modules were stored. A global trigger could also be generated if two or more plastic panels observed coincident events. Hence, LS events could only be stored if signals from at least one crystal or two plastic panels were triggered.
Variables from the laboratory environment and detector stability parameters were monitored. Detectors event rate, behavior of parameters used for event selection, and other detector aspects related to its stability were daily monitored by the collaboration. Environment variables were also controlled through the Grafana application [36], as for example the PMTs voltage levels, relative humidity, radon air concentration and the detectors temperature. More details of the COSINE-100 experiment detectors performance, data acquisition system, and environmental monitoring are reported in Refs. [37, 38, 39, 40].
@@ III DM-electron scattering model
According to momentum and energy conservations, WIMP-like DM deposits more energy in the target when scattering off particles with comparable mass. Hence, the recent focus on searches for light DM, especially below 10 GeV, has motivated the increasing interest in WIMP-like DM interaction with electrons, since in this mass range, DM particles can produce detectable signals above the threshold of current detectors.
In the following subsection III.1, the main concepts and formulas relevant for direct detection searches within the DM-electron scattering model are presented. Subsection III.2 provides a brief discussion of the computation of atomic ionization factors, including the impact of electronic wavefunctions modeling on the expected DM–electron signals and the choice of models appropriate for the analysis presented in this work.
@@ III.1 Overview and key expressions for direct detection
Although the standard DM-nucleon scattering scenario assumes a weak interaction [41], there are no restrictions in the models to DM interacting simultaneously with nucleons and electrons, despite a new vector boson is commonly postulated in the latter framework [18]. The most frequently considered model for the interaction between sub-GeV DM and electrons suggests a hidden sector boson to be the mediator, without restrictions to its spin and parity [31]. However, in order to simplify the model and accommodate the well motivated scenario of the dark photon [16], this boson is supposed to be a vector. In this search, we consider only the scattering between DM and electrons, neglecting the DM-nucleon interaction, in order to directly compare the analysis with reported studies from other experiments. Nevertheless, the typical momentum transfer involved in the scattering is much smaller than the sodium and iodine nuclei masses, implying that the nuclear recoil energy would be negligible. In this scenario, the Feynman diagram representing the scattering is presented in Fig. 2.
Figure 2: Feynman diagram that represents the DM-electron scattering, where $V$ is a vector boson and $\chi$ is the DM particle.
Generally, describing the kinematics of the electron ionization by a DM scattering is more complex than the typical DM-nucleon interaction, because the possible values of the momentum transfer in the interaction ($q$) can be as small as O(100 eV) for DM masses in the MeV range and energy depositions of O(1-10 eV), or as high as O(10 MeV) for DM masses in the GeV range and energy depositions of O(1-10 keV) [42, 31].
Focusing on the direct detection of electron recoils, the velocity averaged cross section for the DM-electron scattering in a detector target can be determined by [20]:
@@
where $m_{e}$ is the electron mass, $f(v_{\chi})$ and $v_{\chi}$ are the DM velocity distribution and DM velocity in the Earth frame, respectively, $a_{0}$ is the Bohr radius, written as $a_{0}=\frac{1}{m_{e}\,\alpha}$, $f_{ion}(E,q)$ is the ionization factor, and $q_{+}$ and $q_{-}$ are the maximum and minimum momentum transfer, respectively, given by $q_{\pm}=m_{\chi}v_{\chi}\pm\sqrt{m^{2}_{\chi}v_{\chi}^{2}-2m_{\chi}E}$, where $m_{\chi}$ is the DM mass, and $E$ is the energy deposition.
The DM form factor ($F_{\chi}(q)$) and the WIMP-electron scattering cross section (${\overline{\sigma}}_{e}$) are the main expressions that contain information about the interaction mediator. These are given by [32]:
$F_{\chi}\,(q)=\frac{\big(\frac{m_{\chi}}{m_{e}}\big)^{2}+\alpha^{2}}{\big(\frac{m_{\chi}}{m_{e}}\big)^{2}+(\alpha\,a_{0}\,q)^{2}}$ (2)
$\overline{\sigma}_{e}=\frac{16\,\pi\,a_{0}^{2}\,\alpha^{2}\,\alpha_{\chi}^{2}}{\Big(\big(\frac{m_{\chi}}{m_{e}}\big)^{2}+\alpha^{2}\Big)^{2}}$ (3)
where $\alpha$ is the fine structure constant, $\alpha_{\chi}$ is the DM-electron coupling strength. $\overline{\sigma}_{e}$ is the interaction cross section considering a free electron whose scattering with the DM particle transferred a momentum of $a^{-1}_{0}$ [43].
Since $\overline{\sigma}_{e}$ and $F_{\chi}(q)$ depend on the mediator mass, studies usually consider two limiting cases, such that a standard scenario is established, and a precise comparison between different results becomes simple. The first case considers a heavy mediator ($m_{\chi}>>\alpha\,m_{e}$), therefore, $F_{\chi}(q)\longrightarrow 1$. The opposite case considers a light mediator ($m_{\chi}<<\alpha\,m_{e}$), therefore, $F_{\chi}(q)\longrightarrow\dfrac{1}{(a_{0}\,q)^{2}}$. / The standard halo model has been assumed in this analysis:
@@
where $\vec{v}$ is the DM velocity in the galactic frame, $\vec{v_{E}}(t)$ is the Earth velocity in the galactic frame, $v_{0}=220$ km/s is the Sun rotation velocity with respect to the galactic center, and $v_{esc}=550$ km/s is the Milky Way escape velocity. $\vec{v_{E}}(t)$ depends on the relative velocity of Earth with respect to the Sun, and the peculiar velocity of the solar system in the galactic frame. This parametrization is in accordance with the procedure described in Ref. [44].
@@
where $n^{NaI}_{a}$ is the atomic number density per kilogram of NaI, and $\rho_{DM}=0.4$ GeV/cm3 is the approximate energy density of CDM at Earth [45, 31].
The parameters chosen for the DM velocity distribution and local density are the same as those adopted in Ref. [32], enabling a direct comparison with the DAMA/LIBRA constraints. It is important to note that these values differ from those typically assumed in DM–electron scattering searches [46]. The constraints presented in this work are expected to be less than 25% more stringent than those obtained using the standard parameter choices. Therefore, comparison with results from other direct detection experiments should be made with appropriate care.